OpenStudy (anonymous):

Hi everyone! How would you go about solving this: y''+x^2y'+e^x=0 Can someone push me in the right direction? Thanks! :o)

OpenStudy (amistre64):

find the homogenous and then the particular solutions

OpenStudy (amistre64):

y'' + x^2 y' = 0 gives us the homogenous solution if we multiply thru by say e^(1/3 x^3) that should help out if memory serves

OpenStudy (anonymous):

what method is that? I can only seem to recall how to solve constant coefficient homogeneous de's

OpenStudy (amistre64):

what would you have multiplied thru if it was a constant there? say 3 e^(3x) right?

OpenStudy (anonymous):

if there was a constant where exactly?

OpenStudy (amistre64):

yousaid you remembered it, insted of x^2 there is a 3

OpenStudy (anonymous):

I wouldn't multiply through at all... I would make the equation r^2+3r+1=0, find the roots and apply the proper formula for the solution

OpenStudy (amistre64):

y' + 3y = 0 let u = e^3x, u' = 3 e^(3x) uy' + 3uy = 0 e^(3x) y' + 3 e^(3x) y = 0 this is the product rule if we integrate both sides e^(3x)y = 1+C right?

OpenStudy (anonymous):

yeah I don't think I have been taught this method yet...looks completely foreign to me

OpenStudy (amistre64):

its taught before the characteristic equation method, since it relies only on what you learnt in derivatives class

OpenStudy (anonymous):

hmm...maybe I just forgot then...so basically flip it around, integrate, and what not?

OpenStudy (amistre64):

(uy)' = uy' + u'y let u = e^(3x), and y=y e^(3x) y' + 3 e^(3x)y = e^(3x) (y' + 3y)

OpenStudy (amistre64):

yeah, determine an integration function i think is the technical name for it, that may turn it into a reverse product rule

OpenStudy (anonymous):

okay...that's good enough...I can look back at my notes...my teacher never had us taking double integrals before though...I will review...Thanks!

OpenStudy (amistre64):

y'' + x^2 y' = -e^x let u = e^(??) and u' = x^2 e^(??), ?? has to be 1/3 x^3 multiply both sides by e^(1/3 x^3) e^(1/3 x^3) y'' + x^2 e^(1/3 x^3) y' = - e^(1/3 x^3) e^x e^(1/3 x^3) y' = -integral [e^(1/3 x^3 +x)] y' = -integral [e^(1/3 x^3 +x)] * e^(-1/3 x^3)

OpenStudy (amistre64):

let v = e^(1/3 x^3 +x) ln(v) = 1/3 x^3 + x int(ln(v)) = 1/12 x^4 + 1/2 x^2 + C v ln(v) - v = 1/12 x^4 + 1/2 x^2 + C there might be a simpler approach, but this is what came to mind to start with :)

OpenStudy (anonymous):

yeah, seeing all that, I think the teacher wanted me to see if a 2nd solution besides zero existed using that big integral formula, so y=0 should be the only solution then I'm thinking

OpenStudy (anonymous):

yeah I definitely haven't solved 2nd degree variable coeff before :o)

OpenStudy (amistre64):

well if your satisified with it, then we are done here :) good luck and all

OpenStudy (anonymous):

thanks for the hard work! I will always be able to come back and see you mathematical wizardry! :o)